# Arrival Time Prediction for Autonomous Shuttle Services in the Real World: Evidence from Five Cities

### v1

@@ VI-C Discussion
Since the data is gathered during pilot tests, the data availability is limited. Most sites provide only GPS positioning and speed data. Given this constraint, we cannot rely on many of the input features that have been referenced in prior literature, such as passenger data, that have been proven to be the most impactful for dwell time prediction [26]. Moreover, it is reasonable to assume that demand is highly volatile during pilot phases, as some passengers might be merely curiosity-driven, keen on experiencing technological advancement just once. In some pilot sites, such as Tampere, the shuttle is skipping the majority of the stops due to limited demand. This variability further complicates dwell time predictions. Given that the dwell time of zero emerged as a challenge in our prediction models and the error in dwell time prediction had more impact on the overall AT prediction, passenger monitoring promises a great influence on the final AT prediction.
@@
Regarding running time prediction, we also encounter challenges. As autonomous shuttles operate on both shared public roads and private lanes or shared lanes only with pedestrians and cyclists, we cannot draw from standard road traffic speed estimates available from car data. Instead, we are limited to the most recent running time observations from the shuttle. In some test scenarios where only a single vehicle was operational, or operations were stopped due to technical issues, these recent observations can quickly become outdated and, thus, less reliable. Further, local regulations constrain the speed limit of the shuttle, limiting the variance in running times and, therefore, the potential improvement obtained by predictions.
@@
Our key recommendation for practitioners collecting data during pilots is to be mindful of data consistency. Specifically, if ordinary operations are interrupted, data collection should either be paused, or the collected data should be labeled as ’non-ordinary’. Without this information, we cannot guarantee that test runs to check technical feasibility without the intention of transporting passengers are not included in the data set. Furthermore, any changes to the settings of the shuttle, such as dwelling settings, the sampling rate of the GPS device, or acceleration/deceleration behavior, should be documented. This would greatly improve data pre-processing and the evaluation of predictions for regular operations post-pilot.
VII Conclusion
This work studied the impact of autonomous vehicles on prediction models within public transport. As autonomous transportation emerges, accurate arrival time predictions are vital to gain public acceptance. Given the early stage of this technology, models typically have to be applied to data originating from pilot tests. Hence, we emphasized the importance of careful data collection to practitioners to obtain an estimate of the prediction for regular operations post-pilot. We tested several established machine learning and deep learning models on dwell and running time prediction in the particular setting of autonomous mobility using data from five diverse pilots across Europe. A key finding is the absence of a universally best model for dwell time prediction and that the performance depends on the specific characteristics of the pilot sites. In cases with high data availability and a long, frequently-used route, the hierarchical approach of a random forest classifier combined with a graph neural network is superior. However, in scenarios where the dwell time is mainly automated and does not significantly vary with passenger density, the prediction challenge simplifies to a classification task, and the tree-based models are preferable. Surprisingly, leveraging deep learning architectures that incorporate spatial information for running time predictions failed to yield clear benefits over the tree-based methods. This observation can be partly attributed to the fact that the shuttles are mainly deployed in low-traffic and low-speed zones or are constrained by a regulatory speed limit for automated shuttles. For the final AT prediction, dwell time prediction proves to be the driving component, with a critical factor of the prediction model being able to detect when shuttles skip stops. Encouragingly, the results for the final AT prediction are promising, with maximum errors not exceeding 60 seconds for predictions up to five stops ahead.
@@ VII Conclusion
Future research directions are threefold. First, as autonomous public transport systems evolve in complexity, research should monitor performance changes in predictions in response to growing passenger demand and potential increases in speed limits. Secondly, there is room to explore other models, such as combining the hierarchical zero-inflated approach into one model with combined optimization, e.g., with a Gaussian mixture model. Lastly, future research should collect and incorporate additional data features into the prediction models. Specifically, passenger data, along with details on shuttle characteristics or operational settings that affect dwell time or acceleration, declaration, and cruising speed, are potentially important data characteristics that influence the prediction of the arrival time of automated shuttles.
Acknowledgements
This project has received funding from the European Union’s Horizon 2020 research and innovation program under grant agreement No 875530.

### v2

@@ V-G Key insights from working with data from real-world automated shuttle deployment
Data gathered during pilot tests often comes with limitations. Most sites provide only GPS positioning and speed data. Given this constraint, we cannot rely on many of the input features that have been referenced in prior literature, such as passenger data, which have been proven to be the most impactful for dwell time prediction [RASHIDI202332]. It is reasonable to assume that demand is highly volatile during pilot phases, as some passengers might be merely curiosity-driven, keen on experiencing technological advancement just once. This variability further complicates dwell time predictions. Given that the dwell time of zero emerged as a challenge in our prediction models and the error in dwell time prediction had more impact on the overall AT prediction, passenger monitoring promises a great influence on the final AT prediction.
@@
Regarding running time prediction, we also encounter challenges. As automated shuttles operate on both shared public roads and private lanes or share lanes only with pedestrians and cyclists, we cannot draw from standard road traffic speed estimates available from car data. Instead, we are limited to the most recent running time observations from the shuttle. In some test scenarios where only a single vehicle was operational, or operations were stopped due to technical issues, these recent observations can quickly become outdated and, thus, less reliable. Further, local regulations constrain the speed limit of the shuttle, limiting the variance in running times and, therefore, the potential improvement obtained by predictions.
@@
In the previous section, we quantified the impact of deviations from scheduled operations that were not documented in the data. Thus, our key recommendation for practitioners collecting data during pilots is to be mindful of data consistency. If ordinary operations are interrupted, data collection should either be paused or the collected data should be labeled as ’non-ordinary’. Without this information, we cannot guarantee that test runs to check technical feasibility without the intention of transporting passengers are not included in the data set. Similarly, given that the vehicles are automated at level 4 with a safety driver who may interfere with operations— such as letting passengers board between stops— it is essential to flag any instances of safety driver intervention. Failing to do so could undermine the validity of conclusions drawn about level 5 automation. Furthermore, any changes to the settings of the shuttle, such as dwelling settings, the sampling rate of the GPS device, or acceleration/deceleration behavior, should be documented. This would greatly improve data processing and the evaluation of predictions for regular operations post-pilot. The conclusions drawn in this study are scoped to the current deployment context of automated shuttles, i.e., smaller services that complement existing public transport. We deliberately do not generalize to high-capacity trunk services on primary roads, where higher speed limits and traffic volumes may change the value of running time prediction, warranting a dedicated study. As technology and deployments evolve, model choices should be re-evaluated.
VI Conclusion
This work studied the impact of automated vehicles on prediction models within public transport. As autonomous transportation emerges, accurate AT predictions are vital to gaining public acceptance. Given the early stage of this technology, models typically have to be applied to data originating from pilot tests with automation level 4 (with a safety driver). Hence, we emphasized the importance of careful data collection to practitioners to obtain an estimate of the prediction for regular operations post-pilot.
We tested several established machine learning and deep learning models on dwell and running time prediction in the setting of automated mobility using data from six diverse pilots across Europe. A key finding is the absence of a universally best model for dwell time prediction. In cases with high data availability, a long, frequently-used route, and deployment in suburban residential areas, the hierarchical approach of a random forest classifier combined with a graph neural network is superior. However, in scenarios where the dwell time is mainly automated, does not significantly vary with passenger density, and the service type has lower passenger counts (e.g., employee-only shuttles), the prediction challenge simplifies to a classification task, and the tree-based models are preferable.
We found that dwell time predictability declines when (i) the pilot site’s focus extends beyond passenger transportation, (ii) passenger cars or vans with manual doors and low-clearance boarding are used instead of shuttle buses, and (iii) data availability is limited. The greatest improvement in predictability occurs with automated dwell times, where a standard door opening time is applied, underlining the critical role of dwell time prediction in accurate AT estimation for automated shuttles. Overall, advanced prediction models showed a clear advantage over simpler approaches like the Mean or Lag model, which failed to capture data variability.
For running time prediction, however, the benefit of advanced prediction models was less clear. Surprisingly, leveraging deep learning architectures that incorporate spatial information failed to yield clear benefits over tree-based methods. This observation can be partly attributed to the fact that the shuttles are mainly deployed in low-speed zones or are constrained by a regulatory speed limit for automated shuttles. Consequently, for automated shuttles at their current stage of development, the Mean or Lag Regressor already provides adequate explainability for most pilot sites when evaluating individual predictions. However, when the goal is to provide accurate predictions well in advance, a prediction model can help reduce the accumulated error.
To summarize, at the current stage of automated shuttle deployment, we observed contrasting trends in dwell and running time predictions, with advanced models proving more valuable for dwell and less critical for running time prediction. Consequently, for the final AT prediction, dwell time prediction proves to be the driving component, with a critical model capability being able to detect skipped stops. Encouragingly, the results for the final AT prediction are promising, with maximum errors of 60 seconds, equaling maximum absolute percentage errors of $15\%$ for predictions up to five stops ahead.
Most importantly, our findings are not consistent with the main focus of most of the literature, which develops sophisticated (GNN) models for running and travel time prediction. In contrast, dwell time prediction —shown to be especially influential for automated shuttles— has received comparatively less attention and typically relies on simpler models. Additionally, in working with noisy real-world data, our work suggests that factors such as careful data collection, thorough data preparation, and the consideration of operational models are more impactful than the final choice of prediction model.
@@ VI Conclusion
Future research directions are threefold. First, as automated public transport systems evolve in complexity, research should monitor performance changes in predictions in response to growing passenger demand and increased speed limits. Secondly, there is room to explore other models, such as combining the hierarchical zero-inflated approach into one model with combined optimization, e.g., with a Gaussian mixture model. Lastly, future research should collect and incorporate additional data features into the prediction models, i.e., on-board sensor data. As such, compiling a consistent POI/land-use layer is a valuable next step, and combining such static demand proxies with future automatic passenger-count data would likely yield the greatest improvement in dwell time prediction. Specifically, passenger data, along with details on shuttle characteristics and operational settings that affect dwell time, acceleration, and cruising speed, are potentially important data characteristics that influence the AT prediction of automated shuttles.
TABLE VII: Deployment statistics by site.
(a) Network geometry. Network density equals to the total length of drivable roads.
Site Service Area (${\mathrm{km}}^{2}$) Network density (km/km2) Intersections
Linköping $0.94$ $75.00$ $76$
Tampere $0.59$ $81.54$ $123$
Les Mureaux $0.31$ $54.79$ $14$
Madrid $0.05$ $43.65$ 99
Graz $0.30$ $55.33$ $29$
Carinthia $0.23$ $35.61$ $22$
(b) POIs and road-type mix. POIs include amenity, shop, leisure, tourism, office; “POIs @200m” are POIs within 200 m of each stop (median [IQR]). Road types are length-weighted shares of top classes.
Site POIs POIs@200 m Top road types (share %)
Linköping $172$ 27 [17–50] service (39), living (20), $0$ residential (16)
Tampere $235$ 19 [13–30] service (65), residential (20), secondary (8)
Les Mureaux $11$ 1 [0–4] service (100)
Madrid 33 0 [0–1] service (100)
Graz $51$ 14 [6–27] service (66), residential (27), primary (5)
Carinthia $107$ 35 [26–46] residential (66), service (19), primary (15)
We provide a context table derived from OpenStreetMap [OpenStreetMap] (Table VII(b)), computed within each service area. For each site, we report service area ($km^{2}$), drivable road density ($km/km^{2}$) as a measure of network density, consolidated intersection count (3‑way+), total POIs, POIs within 200 m per stop (median [IQR]), and the length‑weighted road‑type mix (e.g., service, primary, residential). These proxies characterize deployment context and complement Tables I-II by quantifying network and demand features.  [^2] Because Open Street Map coverage varies by location, values should be interpreted as context descriptors rather than exhaustive counts.
Table VII(b) underlines the deployment context of automated shuttles as smaller services deployed in residential or service (retail/campus/company) areas. Service-dominated sites such as Les Mureaux have sparse intersections and low POI proximity, conditions under which tree-based models perform strongly. By contrast, mixed-traffic settings (e.g., Linköping, Tampere, Carinthia) exhibit richer connectivity and amenity density, and benefit from spatial models once observation counts are moderate.
[Uncaptioned image] Carolin Schmidt is a Postdoctoral Scholar at the Technical University of Denmark in the Machine Learning for Smart Mobility lab, where her research includes learning-based control for autonomous mobility systems and climate change adaptation. She completed her Master’s degree in Electrical Power Engineering and Operations Research with Honors at RWTH Aachen University. As part of her PhD, she was involved in an EU-wide project on the real-world deployment of autonomous shuttles. During her time as a visiting researcher at the Autonomous Systems Lab at Stanford University, she worked on integrating deep reinforcement learning (RL) with established optimization and control methods to create scalable, robust, and safe RL, contributing to applications in large-scale transportation systems, operations research, and robotics.
[Uncaptioned image] Mathias Niemann Tygesen holds a PhD from the Technical University of Denmark, where he specialized in Spatio-Temporal Machine Learning for public transport. Following a PostDoc on Reinforcement Learning for dynamic road pricing, he now works on demand forecasting in the retail fashion industry, applying advanced machine learning to optimize supply chains and inventory management. His work focuses on translating cutting-edge research into practical solutions across transportation and retail sectors.
[Uncaptioned image] Filipe Rodrigues is associate professor at the Technical University of Denmark (DTU) in the Machine Learning for Smart Mobility (MLSM) lab, where his research is primarily focused on machine learning models for understanding and optimizing urban mobility and human behavior. Previously, he was a H.C. Ørsted / Marie-Skłodowska Curie Actions (COFUND) postdoctoral fellow, also at DTU, working on spatio-temporal models of mobility demand with emphasis on modeling uncertainty and the effect of special events. He has published more than 60 articles in leading conferences and journals in both transportation and machine learning. His research interests span machine learning, reinforcement learning, intelligent transportation systems, and urban mobility.